How to Find (r) 〈 10000%-0.5 × 14400% +0.5 × 6400r
Answer #1

10000r=0.5(14400r + 6400r)

10000r = 0.5 (20800r )

(10000/20800)r = 0.5

(0.481)r =0.5

therefore r will be approximate to 1.